For the given complex numbers
u = 9 - i, v = 4/3 - i
Compute the product uv

Answer by checkley77(12844)

u = 9 - i, v = 4/3 - i
uv=(9-i)(4/3-i)
uv=9*4/3-4/3i-9i+1
uv=12-31/4i+1
uv=13+31i/4 ans.